# Segmenting Objectiveness and Task-awareness Unknown Region for Autonomous Driving

### v1

@@ 4.1. Experimental Setup
Outlier Exposure. Following prior works (Nayal et al., 2023; Liu et al., 2023; Tian et al., 2022), we adopt an outlier exposure (OE) strategy to enhance our model’s ability to distinguish inlier and outlier regions. Specifically, we utilize Cityscapes (Cordts et al., 2016) as the inlier dataset and COCO (Lin et al., 2014) as the outlier dataset to generate synthetic training samples. Cityscapes contains 2975 annotated training images with 19 defined inlier semantic categories, which are considered as normal objects in the driving scene. To introduce OOD objects, we leverage the diverse object categories present in COCO different than Cityscapes inliers and blend them into Cityscapes images, thereby simulating out-of-distribution scenarios. This augmentation approach ensures that our model is exposed to a variety of anomalous objects during training, improving its generalization ability for real-world OOD detection tasks.
@@ 4.2. Main Results
Qualitative Results. Figure 3 presents a visual comparison of anomaly segmentation results produced by three representative methods: RPL (Liu et al., 2023), RbA (Nayal et al., 2023), and our proposed method SOTA. In the figure, the first column shows the original input images, while the subsequent columns display the corresponding anomaly maps of various approaches. It can be observed that RPL tends to generate diffused predictions with substantial false positives, particularly along object boundaries and in areas outside the road region. Although RbA improves upon RPL by offering sharper delineation, it still suffers from incomplete segmentation and residual noise. In contrast, SOTA delivers more accurate and comprehensive segmentation of OOD objects while effectively suppressing extraneous pixel predictions beyond the road area. These qualitative findings underscore the advantages of our approach, which unifies attribute-aware feature learning and scene-guided context modeling to achieve a superior balance between precise anomaly localization and noise reduction. Additional qualitative results are provided in the Supplementary Materials.
@@ 4.3. Ablation Studies
SOTA Ablation Analysis. Table 3 presents an ablation study comparing different configurations of our proposed method SOTA. In configuration (a), no additional modules are applied, representing the baseline RbA output. In configuration (b), only the SG-PCA variant using the raw OOD prompt fusion ($\mathbf{\tilde{y}}$) is applied, effectively using RbA’s anomaly score as the mask prompt input to SAM. Configuration (c) incorporates only the SFB along with mask decoder finetuning, thereby directly integrating the enhanced image features with the anomaly cues. In configuration (d), only the SG-PCA variant using the raw OOD prompt fusion ($\mathbf{\tilde{y}}$) is applied with finetuning. Configuration (e) applies SG-PCA with both the raw OOD prompt fusion ($\mathbf{\tilde{y}}$) and the road semantic prompt ($\mathbf{y^{\prime}_{t}}$); here, the dual cues help refine the spatial context, yielding a distinct performance pattern. Configuration (f) employs SG-PCA with the combination of the raw OOD prompt and the road-related OOD prompt fusion ($\mathbf{\tilde{y}_{t}}$) with finetuning, demonstrating further improvement by incorporating contextual information that accounts for partial object extensions. In configuration (g), both SFB and SG-PCA (using $\mathbf{\tilde{y}}$ and $\mathbf{\tilde{y}_{t}}$) are integrated without finetuning the mask decoder, whereas configuration (h) activates both modules together with finetuning. Notably, configuration (h) achieves the highest performance, whidch yields a relative AuPRC gain and the lowest FPR on both RoadAnomaly and Fishyscapes Lost & Found. This indicates that the combined use of SFB and the SG-PCA variant employing both $\mathbf{\tilde{y}}$ and $\mathbf{\tilde{y}_{t}}$ leads to a more effective integration of anomaly cues and contextual scene information. This analysis suggests that while the SFB enhances the raw anomaly features, the careful selection and fusion of SG-PCA variants via channel-wise attention is critical for suppressing spurious detections and achieving robust anomaly segmentation.
Table 4. Generalization of SOTA to different anomaly scores. We evaluate SOTA by applying it to other detectors, SOTA consistently improves AuROC and AuPRC while reducing FPR, demonstrating its strong generalization ability and compatibility with existing pixel-level anomaly detectors.
@@
Abaltions on Qualitative Results. As corroborated by our qualitative analysis , the visual results reinforce the quantitative improvements observed in Table 3. In Figure 4, column1 and cloumn6 present the original input images and ground truth, while the column2 shows the anomaly outputs produced by RbA. Column3 displays the results when only the SFB is added, revealing enhanced feature disentanglement and more complete anomaly delineation. Column4 presents the outputs obtained by solely incorporating the SG-PCA module, where the channel-wise cross attention effectively suppresses false positives outside the drivable area. Finally, column 5 corresponding to the full SOTA configuration with both modules and mask decoder finetuning—demonstrates the most accurate and robust anomaly segmentation, effectively balancing high precision in OOD detection with minimal background noise. This progressive improvement from RbA to the full SOTA setup underscores the efficacy of our approach in integrating refined anomaly cues with contextual scene information.
@@
Generalize to other anomaly scores. To further illustrate our method’s generalization capability, we integrate the anomaly scores and in-distribution logits from two distinct baselines—RPL and Mask2Anomaly—into our SOTA framework. As shown in Table 4, for the RPL baseline on the RoadAnomaly dataset, our integration leads to a relative improvement of approximately 6.26% in AuPRC and a reduction of about 2.62% in FPR. On the Fishyscapes Lost & Found dataset, our method enhances AuPRC by around 3.71% with a slight decrease in FPR. Similarly, when applied to Mask2Anomaly, SOTA yields substantial AuPRC gains and significant FPR reductions on RoadAnomaly, while maintaining consistent performance improvements on Fishyscapes Lost & Found. These results confirm that our framework is capable of effectively leveraging both anomaly confidence and semantic segmentation cues to robustly enhance OOD detection performance across different baselines.

### v2

@@ 4.1. Experimental Setup
Outlier Exposure. Following prior works (Nayal et al., 2023; Liu et al., 2023; Tian et al., 2022), we adopt an OE strategy to enrich the training distribution. Cityscapes (Cordts et al., 2016) images are used only as unlabeled backgrounds, without exploiting their semantic annotations. Object instances from COCO (Lin et al., 2014), extracted with their ground-truth masks, are composited into Cityscapes scenes to simulate anomalous appearances. Although COCO supervision is used to obtain object regions, no inlier supervision from Cityscapes is involved; hence, the conventional IND–OOD supervision paradigm is not explicitly applied. Instead, COCO simply provides additional object diversity, enabling the model to generalize better to unexpected visual patterns in real-world driving scenarios.
@@ 4.2. Main Results
Qualitative Results. Figure 3 presents a visual comparison of anomaly segmentation results produced by three representative methods: RPL (Liu et al., 2023), RbA (Nayal et al., 2023), and our proposed method SOTA. In the figure, the first column shows the original input images, while the subsequent columns display the corresponding anomaly maps of various approaches. It can be observed that RPL tends to generate diffused predictions with substantial false positives, particularly along object boundaries and in areas outside the road region. Although RbA improves upon RPL by offering sharper delineation, it still suffers from incomplete segmentation and residual noise. In contrast, SOTA delivers more accurate and comprehensive segmentation of OOD objects while effectively suppressing extraneous pixel predictions beyond the road area. These qualitative findings underscore the advantages of our approach, which unifies attribute-aware feature learning and scene-guided context modeling to achieve a superior balance between precise anomaly localization and noise reduction.
@@ 4.3. Ablation Studies
SOTA Ablation Analysis. Table 3 presents an ablation study comparing different configurations of our proposed method SOTA. In configuration (a), no additional modules are applied, representing the baseline RbA output. In configuration (b), only the SG-PCA variant using the raw OOD prompt fusion ($\mathbf{\tilde{y}}$) is applied, effectively using RbA’s anomaly score as the mask prompt input to SAM. Configuration (c) incorporates only the SFB along with mask decoder finetuning, thereby directly integrating the enhanced image features with the anomaly cues. In configuration (d), only the SG-PCA variant using the raw OOD prompt fusion ($\mathbf{\tilde{y}}$) is applied with finetuning. Configuration (e) applies SG-PCA with both the raw OOD prompt fusion ($\mathbf{\tilde{y}}$) and the road semantic prompt ($\mathbf{y^{\prime}_{t}}$); here, the dual cues help refine the spatial context, yielding a distinct performance pattern. Configuration (f) employs SG-PCA with the combination of the raw OOD prompt and the road-related OOD prompt fusion ($\mathbf{\tilde{y}_{t}}$) with finetuning, demonstrating further improvement by incorporating contextual information that accounts for partial object extensions. In configuration (g), both SFB and SG-PCA (using $\mathbf{\tilde{y}}$ and $\mathbf{\tilde{y}_{t}}$) are integrated without finetuning the mask decoder, whereas configuration (h) activates both modules together with finetuning. Notably, configuration (h) achieves the highest performance, which yields a relative AuPRC gain and the lowest FPR on both RoadAnomaly and Fishyscapes Lost & Found. This indicates that the combined use of SFB and the SG-PCA variant employing both $\mathbf{\tilde{y}}$ and $\mathbf{\tilde{y}_{t}}$ leads to a more effective integration of anomaly cues and contextual scene information. This analysis suggests that while the SFB enhances the raw anomaly features, the careful selection and fusion of SG-PCA variants via cross attention is critical for suppressing spurious detections and achieving robust anomaly segmentation.
Table 4. Generalization of SOTA to different anomaly scores. We evaluate SOTA by applying it to other detectors, SOTA shows its strong generalization ability and compatibility with existing pixel-level anomaly detectors.
@@
Ablations on Qualitative Results. As corroborated by our qualitative analysis , the visual results reinforce the quantitative improvements observed in Table 3. In Figure 4, column 1 and column 6 present the original input images and the ground truth, respectively. Column 2 shows the anomaly outputs produced by RbA. Column 3 displays the results when only the SFB module is added, revealing enhanced feature disentanglement and more complete anomaly delineation. Column 4 presents the outputs obtained by solely incorporating the SG-PCA module, where the cross-attention effectively suppresses false positives outside the drivable area. Finally, column 5, which corresponds to the full SOTA configuration incorporating both modules and mask decoder fine-tuning, achieves the most accurate and robust anomaly segmentation, effectively balancing high precision in OoD detection and minimal background noise. This progressive improvement from RbA to the full SOTA setup underscores the efficacy of our approach in integrating refined anomaly cues with contextual scene information.
@@
Generalize to other anomaly scores. To further illustrate our method’s generalization capability, we integrate the anomaly scores and in-distribution logits from two distinct baselines, RPL and Mask2Anomaly, into our SOTA framework. As shown in Table 4, for the RPL baseline on the RoadAnomaly dataset, our integration leads to a relative improvement of approximately 6.26% in AuPRC and a reduction of about 2.62% in FPR. On the Fishyscapes Lost & Found dataset, our method enhances AuPRC by around 3.71% with a slight decrease in FPR. Similarly, when applied to Mask2Anomaly, SOTA yields substantial AuPRC gains and significant FPR reductions on RoadAnomaly and on Fishyscapes Lost & Found. These results confirm that our framework is capable of effectively leveraging both anomaly confidence and semantic segmentation cues to robustly enhance OOD detection performance across different baselines.
@@ 5. Conclusions
6. Acknowledgements
This work was supported by the National Key Research and Development Program of China under Grant No. 2023YFA1008500. Guanglei Yang is supported by the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20233458 and the Postdoctoral Fund of Harbin Institute of Technology under Grant Number AUGA5630115523.
Appendix
Appendix A Implementation Details
A.1. Task-aware Aggregation Module
To extract a comprehensive representation of the drivable area, the Task-aware Aggregation Module first identifies the road class from the semantic segmentation output and generates a binary road mask. To account for perspective distortions and partial occlusions, the mask is refined using a dilation process. Specifically, a $15\times 15$ convolutional kernel filled with ones is applied iteratively 15 times. In each iteration, a padded convolution is performed followed by thresholding 0 to maintain a binary format. This operation results in an expanded road prompt $\mathbf{y}_{t}$ that captures uncertain boundary regions, ensuring that out-of-distribution objects physically located on the road are not excluded from further processing.
A.2. Projection and Alignment Module
Projection Module. The projection module is designed as a lightweight learnable convolutional sub-network that transforms a raw OoD confidence map into an intermediate semantic representation. Given a single-channel input of size $1\times 256\times 256$, it is processed through two $2\times 2$ convolutional layers with stride 2, reducing the spatial resolution to $64\times 64$ while expanding the channel dimension to 16. Each layer is followed by LayerNorm and GELU activation. The resulting feature map has a shape of $16\times 64\times 64$.
Alignment Module. To align the intermediate semantic features with the image encoder’s representation, we apply a final $1\times 1$ convolutional layer that maps the 16-channel feature map to a 256-dimensional embedding. This operation preserves spatial structure and outputs a tensor of shape $256\times 64\times 64$, which is used for downstream fusion with image features.
Table 6. Pixel-level results on Road Anomaly (Lis et al., 2019) and Fishyscapes L&F (Blum et al., 2021). We show the best results in bold. Our method is based on the anomaly score from RbA. SOTA notably improves the results in all metrics on both datasets.
Method Road Anomaly Fishyscapes L&F
AuPRC $\uparrow$ FPR $\downarrow$ AuPRC $\uparrow$ FPR $\downarrow$
EAM (Grcic et al., 2023) 69.4 7.7 81.50 4.20
UNO (Delić et al., 2024) 88.50 7.4 81.80 1.30
RbA (Nayal et al., 2023) 90.28 4.92 80.35 4.58
SOTA (Ours) 92.77 3.48 83.93 2.23
A.3. Outlier Exposure
During the preparation of the training dataset, we adopt the Outlier Exposure (OE) strategy as described in (Nayal et al., 2023; Liu et al., 2023). Specifically, we use the Cityscapes dataset as the inlier background,which comprises 2,975 training images representing typical urban driving scenes. Then we select objects from the COCO dataset as OoD objects after excluding those categories that overlap with the Cityscapes inlier classes. To simulate realistic anomaly scenarios, we paste the selected COCO objects onto the Cityscapes images with a probability of 0.9 for each image, thereby ensuring a high density of anomalous instances during training. Subsequently, we apply the RbA framework (Nayal et al., 2023) to these composite images to compute both anomaly scores and semantic logits, resulting in a comprehensive training dataset that robustly supports OoD segmentation.
Appendix B Additional Results
B.1. Results with Extra Data
Recent methods have pursued enhanced outlier exposure (OE) through diversified data augmentation. UNO (Delić et al., 2024) and EAM (Grcic et al., 2023) expand OE by introducing dual-source augmentation: they incorporate Vistas (Neuhold et al., 2017) as additional in-distribution (ID) data to cover diverse urban layouts (e.g., tunnels, rural roads) while simultaneously leveraging ADE20k (Zhou et al., 2017) as an extended out-of-distribution (OOD) source, thereby increasing OOD category coverage from COCO’s (Lin et al., 2014) 80 classes to over 150. This strategy aims to improve generalization by maximizing data variety, yet risks diluting anomaly discriminability through overlapping distributions and increased annotation complexity.Inspired by their ID expansion but diverging in OOD strategy, we adopt a focused OE paradigm: augmenting ID data with Vistas (combined with Cityscapes (Cordts et al., 2016)) while retaining COCO as the exclusive OOD source.
As shown in Table 6, with the extra data , SOTA consistently outperforms the baselines on both the Road Anomaly and Fishyscapes Lost & Found benchmarks. On Road Anomaly, compared to the baseline RbA, SOTA achieves a significant relative improvement in AuPRC along with a substantial reduction in the FPR.Moreover, SOTA performance far exceeds that of other methods. And On Fishyscapes Lost & Found, although UNO exhibits a slightly lower FPR, our approach yields a markedly higher AuPRC relative to both RbA and UNO. These relative improvements demonstrate that our method delivers more robust anomaly segmentation performance across diverse datasets, even when using additional data.
Table 7. Ablation study on fusion strategy.
Fusion RoadAnomaly Fishyscapes L&F
AuPRC $\uparrow$ FPR $\downarrow$ AuPRC $\uparrow$ FPR $\downarrow$
Add 91.57 4.43 66.80 3.94
SFB 92.46 4.03 76.10 3.53
B.2. Semantic Fusion Block Efficiency Analysis
We conduct an ablation study to evaluate the effectiveness of the proposed Semantic Fusion Block(SFB) module by comparing it with a naive feature fusion strategy, where the OoD-derived features are directly added to the image features. As shown in Table 9, the SFB achieves a relative AuPRC improvement of 1.0% on RoadAnomaly and 9.3% on Fishyscapes Lost&Found, along with consistent reductions in FPR. This demonstrates that our task-aware fusion mechanism more effectively captures semantic inconsistencies by dynamically balancing anomaly and visual cues, leading to enhanced segmentation performance, especially in complex urban environments.
B.3. Abaltion on Finetune Stategies
Table 8. Ablation under different finetuning strategies.
LoRA FFT RoadAnomaly Fishyscapes L&F
CA SA MLP AuPRC $\uparrow$ FPR $\downarrow$ AuPRC $\uparrow$ FPR $\downarrow$
92.49 3.97 73.48 5.46
92.09 5.09 70.09 5.44
92.46 4.03 76.10 3.53
91.91 4.75 71.87 4.88
As detailed in Table 8, our analysis of parameter-efficient finetuning strategies reveals critical trade-offs between domain specialization and cross-dataset generalization. FFT stands for fine-tuning the parameters of the entire mask decoder, achieving impressive performance. Selectively adapting only cross-attention layers (CA LoRA) results in a measurable decline in AuPRC (approximately 0.4% reduction) and a 1.1% increase in FPR compared to FFT, highlighting the limitations of isolated module adaptation. Strikingly, the combined adaptation of cross-attention and MLP layers (CA+MLP LoRA) retains 99.9% of FFT’s accuracy on RoadAnomaly while improving cross-dataset generalization on Fishyscapes L&F by 2.72%, with FPR reduced by 1.93%—a synergy attributed to MLP-enabled channel interaction modeling. Conversely, full-module adaptation (CA+SA+MLP LoRA) degrades AuPRC by 0.6% and increases FPR by 0.78% compared to FFT, indicating that excessive parameterization introduces redundancy. These results collectively emphasize that strategic partial adaptation—optimizing only 40% of trainable parameters in CA+MLP LoRA—achieves 97% of FFT’s domain-specific performance while tripling cross-dataset robustness, thereby redefining the efficiency-effectiveness frontier for anomaly segmentation in resource-constrained scenarios.
Table 9. Ablation study with and without dilate refinement.
Dilation & Erosion RoadAnomaly Fishyscapes L&F
AuPRC $\uparrow$ FPR $\downarrow$ AuPRC $\uparrow$ FPR $\downarrow$
92.30 4.13 75.31 3.72
92.46 4.03 76.10 3.53
B.4. Abaltion on Task-aware Aggregation Module
To validate the effect of morphological refinement in the Task-aware Aggregation module, we perform an ablation study by removing the Dilation & Erosion operation applied to the road mask. As shown in Table 9, the inclusion of dilation leads to a relative improvement of 0.16% in AuPRC on RoadAnomaly and 0.79% on Fishyscapes Lost&Found. Additionally, it slightly reduces the FPR on both benchmarks. These results demonstrate that expanding the road prompt helps preserve OoD instances that may partially overlap with the road, thus mitigating segmentation errors caused by limited spatial coverage.
B.5. Qualitative Anomaly Segmentation.
As shown in Fig. 6 , Fig. 7 and Fig. 8, our method SOTA demonstrates superior anomaly segmentation aross multi datasets compared to other mthods. The visual results show that while RPL (Liu et al., 2023) and RbA (Nayal et al., 2023) produce incomplete segmentation and exhibit false positives along object boundaries, SOTA delivers complete and coherent segmentation masks. In particular, our method accurately delineates out-of-distribution objects while effectively incorporating scene context, ensuring that anomalies are precisely localized and reducing spurious detections in non-critical areas. These qualitative improvements confirm that by fully segmenting anomalies and integrating contextual scene information, SOTA offers enhanced robustness and precision in real-world settings.
Appendix C Failure Cases
Despite the robust performance of our method, certain failure cases remain. As illustrated in Fig. 9, when the OoD objects are very small, the pixel-wise OoD detector may fail to register these regions. This results in an absence or near-zero values in the OoD confidence map. In such cases, the subsequent steps in our pipeline cannot recover the missing anomaly cues, leading to missed detections. For instance, as shown in the second column of Fig. 9, RbA does not identify these diminutive objects effectively, and the derived anomaly scores do not capture their presence, which in turn causes our method to struggle in detecting them. This limitation highlights the dependency of our approach on the initial quality of the pixel-wise anomaly predictions.
Figure 6. Qualitative Results on RoadAnomaly dataset.
Figure 7. Qualitative Results on Fishyscapes Lost&Found dataset.
Figure 8. Qualitative Results on SMIYC dataset.
Figure 9. Failure cases.